OpenStudy (anonymous):

Polar equation to rectangular equation, help please? r=2asin(theta)+2acos(theta) I know that r=sqrt(x^2+y^2), but I can't figure out what to do after that point.

OpenStudy (anonymous):

\[x=r \cos \theta,\cos \theta=\frac{ x }{ r },y=r \sin \theta,\sin \theta=\frac{ y }{ r}\] substitute and solve.

OpenStudy (anonymous):

you followed

OpenStudy (anonymous):

Not really... I did this: sqrt(x^2+y^2)=2asin(theta)+2acos(theta) x^2+y^2=2a^2sin(theta)+2a^2cos(theta) x^2+y^2=2a^2(y/r)+2a^2(x/r) But that seems wrong and I don't know what to do after that?

OpenStudy (anonymous):

\[r=\frac{ 2ay }{ r }+\frac{ 2ax }{r },r^2=2a \left( y+x \right)\]

OpenStudy (anonymous):

you got it.

OpenStudy (anonymous):

My answer key says it should be (x-a)^2+(y-a)^2=2a^2 ?

OpenStudy (anonymous):

x^2+y^2-2ax-2ay=0 complete the squares.

OpenStudy (anonymous):

Ohh okay! Thank you so much for your help! :)

OpenStudy (anonymous):

yw